5. (3) B acid is : 3. (12) Complete the following table: [H3O+] [OH-] pH рон Acidic, Basic, Neutral 6.4 x 10-5 1.73 9.02

need help solving this
Answer #1

Гно]Сон] = [ои ) = 1.0 x lo 1.0 x 16 - 1.0 x 16 СИ,0* 6:ч х/б Now Cou] - о. 16 , lo | Сон] = 1.6 xll pH = -log (H₂0]IPOH = 9.81] pH < 14, thus, solution Since is lacidic ] pH - 1.73 Now PH = 14-1.73 pOH = 12.27 Now pH = -log CH₂ot] - log [Hz(pH = 4.90 Now Now PH = -log [Hz ot] & log (Hz0+] = 4.90 or log (Hz ot] = -4.90 → [420] - antilog (-4.98) & [H₃O+] = 1.05 xlo